OpenStudy (anonymous):

Simplify 0.5(√K/√L) divided by 0.5(√L/√K).

OpenStudy (anonymous):

[0.5 * k^0.5 * L^-0.5]/[0.5 * L^0.5 * K^-0.5] 0.5/0.5 = 1 the negative exponents switch top/bottom, thus it simplifies to k/l

OpenStudy (anonymous):

\[\left(\left(\frac{1}{2}\right)\left(\frac{\sqrt{K}}{\sqrt{L}}\right)\right) /\left( \left(\frac{1}{2}\right)\left(\frac{\sqrt{L}}{\sqrt{K}}\right)\right) \]\[\left(\frac{\sqrt{K}}{2 \sqrt{L}}\right) /\left( \frac{\sqrt{L}}{2 \sqrt{K}}\right) \]\[\frac{K}{L} \]